# ADD 2023: Towards Audio Deepfake Detection and Analysis in the Wild

### v1

@@ Abstract
The growing prominence of the field of audio deepfake detection is driven by its wide range of applications, notably in protecting the public from potential fraud and other malicious activities, prompting the need for greater attention and research in this area. The ADD 2023 challenge goes beyond binary real/fake classification by emulating real-world scenarios, such as the identification of manipulated intervals in partially fake audio and determining the source responsible for generating any fake audio, both with real-life implications, notably in audio forensics, law enforcement, and construction of reliable and trustworthy evidence. To further foster research in this area, in this article, we describe the dataset that was used in the fake game, manipulation region location and deepfake algorithm recognition tracks of the challenge. We also focus on the analysis of the technical methodologies by the top-performing participants in each task and note the commonalities and differences in their approaches. Finally, we discuss the current technical limitations as identified through the technical analysis, and provide a roadmap for future research directions. The dataset is available for download [^1] http://addchallenge.cn/downloadADD2023.
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Recent rapid advancements in text-to-speech (TTS) [1] and voice conversion (VC) [2] technologies over the past decades have made it possible to generate high-quality and realistic audio that can be difficult to distinguish from real audio with the naked ear. Such a technology has a potential to be abused and misused, notably in generating deepfake audio for impersonation, fraud, and other malicious purposes. The rapid progress in speech synthesis and voice conversion technologies means that these attacks can be readily launched by anyone with a computer and a microphone, and rapidly spread through social media. Therefore, it is urgent to develop effective deepfake audio detection methods to protect the public from being deceived by deepfake audio.
@@ I Introduction
In response to this growing need to advance deepfake audio detection research in this direction, we launched a second Audio Deepfake Detection Challenge (ADD 2023) [^2] http://addchallenge.cn/add2023 to further promote research on deepfake audio detection and analysis. The tasks in the challenge are designed to emulate real-world scenarios and to motivate research that goes beyond the traditional binary classification of real and fake audio. To this end, we also release a dataset for the ADD 2023 challenge, partitioned into four subsets, each corresponding to a task in the challenge. The dataset is described in detail in Section IV. In addition, we analyze the top-performing participating systems in each track and discuss the future directions of deepfake audio detection research in Section V and Section VI, respectively.
The contributions of this article include: (1) we detail the datasets used in all tracks of the ADD 2023 challenge and provide a comprehensive description of each of the datasets; (2) we analyze the technical methodologies used by the top-performing participants in each task of the ADD 2023 challenge and note the commonalities and differences in their approaches and (3) we discuss the current technical limitations identified during the challenge and provide a roadmap for future research directions. We hope that the insight gained from our analysis will help to further advance the research on deepfake audio detection and analysis, and that the ADD 2023 challenge will serve as a stepping stone for future research in this area.
@@ III Evaluation Metrics
III-A Track 1.1: Generation task (FG-G)
@@
III-B Track 1.2: Detection task (FG-D)
@@
III-C Track 2: Manipulation region location (RL)
The evaluation metric for the RL track is the sentence-level accuracy ($A_{s}$) and the frame-level $\text{F}_{1}$ score ($F_{1f}$). The overall performance of each team in the RL track (Score) is thus evaluated as a weighted sum of $A_{s}$ and the $F_{1f}$, with more emphasis on the latter, which is more challenging to achieve and also the main objective of this track, without ignoring the overall $A_{s}$, which is also important in real-world applications.
III-D Track 3: Deepfake algorithm recognition (AR)
@@
IV Datasets
In this section, we present and describe the datasets [^3] Available for download at http://addchallenge.cn/downloadADD2023 used in the ADD 2023 challenge. Each dataset is used for a specific task that corresponds to a track in the challenge.
IV-A Track 1.1: Generation task (FG-G)
The FG-G track being focused on the generation of deepfake audio, its dataset is composed of speech corpora with transcripts. In this case, we use the AISHELL-3 dataset [8] in the training process, which is a large-scale multi-speaker Mandarin speech dataset for TTS that contains roughly 85 hours of speech data from 218 native Mandarin speakers, totalling 88,035 utterances, each with corresponding transcripts in both Chinese characters and Pinyin.
The testing set for the FG-G subtrack, summarized in Table I, is composed of 998 text sentences (499 per each round) in Simplified Chinese characters, sourced from the Internet, as well as four speaker IDs (two per each round) from the aforementioned AISHELL-3 dataset. For round 1, speakers with the IDs SSB0539 and SSB1448 were chosen, while for round 2, speakers with the IDs SSB0737 and SSB0817 were selected. The text data is used for the generation task, where participants are tasked with generating deepfake audio that convincingly emulates the characteristics of genuine utterances. Without punctuations, the average length of these sentences is 66.98 characters, with a maximum length of 147 characters, a minimum length of 39 characters, and a variance of 136.78 characters.
@@ IV Datasets
IV-B Track 1.2: Detection task (FG-D)
IV-B1 Training and development sets
In the training and development sets, only the AISHELL-3 dataset is used for genuine speech, totalling 60 speakers and 3,012 utterances for training and 60 speakers and 2,307 utterances for development (see Table II(a)). For fake audio, we use samples generated from the following methods:
IV-B2 Testing set
For the testing set of FG-D dataset, we include audio samples from a wider range of sources for both real and fake audio (see Table II(b)). The genuine audio portion comes from the following sources:
  • Tencent121212hhttps://www.tencentcloud.com/products/tts: A TTS and VC platform developed by Tencent providing high quality speech synthesis services.
IV-C Track 2: Manipulation region location (RL)
@@
IV-C1 Training and development sets
The training and development sets (see Table III(a)) contain genuine audio sourced from the training and development sets of the AISHELL-3 dataset and partially fake audio sourced from the training and development sets of the HAD dataset [9].
IV-C2 Testing set
The testing set of Track 2 (see Table III(b)) contains 1,030 speakers and 20,000 utterances for genuine audio from various sources like AISHELL-1, AISHELL-3, THCHS-30, Mobvoi, AliMeeting and Phone recordings. The fake audio samples in the testing set are AISHELL-1 and AISHELL-2 samples spliced with either same-speaker recordings or fake audio generated with the Tacotron 2–LPCNet pipeline, with one or two spliced regions per sentence, as well as utterances processed by the Speech Edit model based on CampNet [26]. The fake audio samples were generated using the same speakers as the genuine audio samples. Genuine utterances are then spliced by replacing certain segments with fake audio or other recordings of the same speaker, following a similar procedure as in [9]. These partially-fake samples total 30,000 utterances from 148 speakers.
IV-D Track 3: Deepfake algorithm recognition (AR)
The objective of the AR track is to recognize the deepfake algorithm used to generate a given piece of audio; therefore its dataset contains genuine utterances as well as fake utterances generated by various deepfake algorithms and commercial TTS platforms (see Table IV).
The dataset used in this track is based on the datasets presented in [29, 30], composed of genuine utterances from the AISHELL-3 dataset and fake utterances generated by various deepfake algorithms from the AISHELL-1 and AISHELL-3 text corpora. The deepfake algorithms used to generate the fake utterances include vocoders like HiFiGAN [10], WaveNet [15] and World [16], as well as commercial solutions from Aliyun [^14] https://ai.aliyun.com/nls/tts, DataBaker [^15] https://www.data-baker.com/specs/compose/online, and AISpeech [^16] https://beta.duiopen.com/openSource/technology/tts.
@@ IV-D Track 3: Deepfake algorithm recognition (AR)
Further augmentations were made to the testing set to emulate real-world scenarios. Initially, a series of high-quality clean audio samples were generated with each of the aforementioned TTS tools. To simulate diverse acoustic environments of the real world, various types and levels of background noise were added to clean audio samples, creating the “noisy” condition to test deepfake algorithm recognition under suboptimal conditions. Additionally, to reflect audio compression commonly found on social media, the Jinshi Video Assistant software [^18] https://www.drmfab.cn/zhushou/ was used to compress clean audio, forming the “compressed” condition. These steps assess the algorithm’s adaptability to encoding changes.
TABLE V: DSR (%) and methods of top-performing systems in Track 1.1 (FG-G) submissions. [“Aug.” = augmentation; “Rep.” = representation; “Arch.” = architecture; “AR” = auto-regressive]
@@
A01 [31] 44.97 Noise + reverb Phoneme seq. Mel spec Mel spec Tacotron 2 WaveRNN - BiLSTM embedding AR
A02 [32] 43.63 - Phoneme seq. Mel spec Mel spec FastPitch HiFiGAN MFA One-hot embedding NAR
A03 [33] 41.48 - Context encoder Linear spec Waveform Hier-TTS - - One-hot embedding AR
A05 [34] 37.35 Concatenation G2P+BERT Mel spec Mel spec FastSpeech 2 HiFiGAN Kaldi One-hot embedding NAR
@@
A06 [35] 30.69 - Phoneme seq. Linear spec Waveform VITS - MAS One-hot embedding NAR
@@
V Technical Analysis
We analyze the technical details of the top-5 performing teams of each track in this section. In the tables are presented not only the performances of top-performing participating teams, but also the results of the baseline systems [6] (denoted in the form of S#, where # is the baseline number).
V-A Track 1.1: Generation task (FG-G)
@@ V Technical Analysis
    Data augmentation: With the policy prohibiting usage of external speech corpora, data augmentation can be worthwhile. Interestingly, among the participating teams of the FG-G task, the use of data augmentation is not widespread, with only A01 [31] and A05 [34] reporting any form of data augmentation. Of those two teams, A01 uses reverberation and additive noises to augment their dataset, while A05 concatenates and re-splices training data belonging to the same speaker ID.
@@
    Text representation: Most teams opted for phoneme sequence as the text representation, which is a common choice for TTS systems, as it is seen to be closer to the phonetic representation of speech, and thus more suitable for speech synthesis, as well as being more robust to spelling errors and out-of-vocabulary words. A notable exception to this trend is Team A03 [33], which used a context encoder to encode the input text into a fixed-length vector. Another deviation from this is Team A05 [34], who used a grapheme-to-phoneme model to convert the input text into phoneme sequence, and used BERT for disambiguating homographs.
@@
    Architectures: The use of non-autoregressive models like FastSpeech 2 is common in this track, which allows for variations in generation pipelines in the form of duration models. Aside from using models that are non-autoregressive, most teams adopt approaches with intermediate audio representations like Mel spectrogram, which are then passed through vocoder models to generate the final audio waveform. It is worth noting, however, that A06 [35] used the VITS architecture, which is fully end-to-end, with its own alignment module of Monotonic Alignment Search, that allows for the generation of audio waveform directly from phoneme sequence by using a variational autoencoder-like encoder-decoder architecture to directly decode into audio waveform.
TABLE VII: Scores (%), sentence accuracies $A_{s}$ (%), frame-wise F1 scores (%) and methods of top-performing systems and baseline in Track 2 (RL). [“Aug.” = augmentation]
Team Score ($\uparrow$) $A_{s}$ ($\uparrow$) $F_{1f}$ ($\uparrow$) Data Aug. Acoustic Features Back-end Classifiers Model Fusion
C01 [42] 67.13 82.23 60.66 - Wav2Vec 2.0 ResNet–Transformer–LSTM -
@@
C05 [46] 59.12 74.52 52.53 Noise Raw, Wav2Vec 2.0 AASIST, FC layer Weighted frame-wise
@@
V-B Track 1.2: Detection task (FG-D)
@@
    Data augmentation: The use of data augmentation in Track 1.2 is widespread and has found much success, with most teams opting to use MUSAN [52] noise augmentation. The use of reverberation is also popular, with most teams opting to add reverberation to their training data to improve the robustness of their models. Interestingly, Team B01 [36] opted to furthermore use copy synthesis as one of their data augmentation methods to obtain more training data from the AISHELL-3 corpus, and given the success of their system, this approach may have been effective in the task.
@@
V-C Track 2: Manipulation region location (RL)

### v2

@@ Abstract
The growing prominence of the field of audio deepfake detection is driven by its wide range of applications, notably in protecting the public from potential fraud and other malicious activities, prompting the need for greater attention and research in this area. The ADD 2023 challenge goes beyond binary real/fake classification by emulating real-world scenarios, such as the identification of manipulated intervals in partially fake audio and determining the source responsible for generating any fake audio, both with real-life implications, notably in audio forensics, law enforcement, and construction of reliable and trustworthy evidence. To further foster research in this area, in this article, we describe the dataset that was used in the fake game, manipulation region location and deepfake algorithm recognition tracks of the challenge. We also focus on the analysis of the technical methodologies by the top-performing participants in each task and note the commonalities and differences in their approaches. Finally, we discuss the current technical limitations as identified through the technical analysis, and provide a roadmap for future research directions. The dataset is available for download at http://addchallenge.cn/downloadADD2023
@@
1 Introduction
Recent rapid advancements in text-to-speech (TTS) [1] and voice conversion (VC) [2] technologies over the past decades have made it possible to generate high-quality and realistic audio that can be difficult to distinguish from real audio with the naked ear. Such a technology has a potential to be abused and misused, notably in generating deepfake audio for impersonation, fraud, and other malicious purposes. The rapid progress in TTS and VC technologies means that these attacks can be readily launched by anyone with a computer and a microphone, and rapidly spread through social media. Therefore, it is urgent to develop effective deepfake audio detection methods to protect the public from being deceived by deepfake audio.
@@ 1 Introduction
In response to this growing need to advance deepfake audio detection research in this direction, we launched a second Audio Deepfake Detection Challenge (ADD 2023) [^1] http://addchallenge.cn/add2023 to further promote research on deepfake audio detection and analysis. The tasks in the ADD 2023 challenge as well as the data used in each task are designed to emulate real-world scenarios and to motivate research that goes beyond the traditional binary classification of real and fake audio and to further accelerate and foster research on detecting and analysing deepfake audio. We hoped that the insight gained from our analysis will help further advance the research on deepfake audio detection and analysis, and that the ADD 2023 challenge will serve as a stepping stone for future research in this area.
With the successful conclusion of the challenge, in order to further prompt relevant research in audio deepfake detection, we release a dataset for the ADD 2023 challenge, partitioned into four subsets, each corresponding to a task in the challenge. We break down the dataset into four tasks and provide a detailed description of each dataset in Section 4. This dataset is designed to simulate different acoustic environments, adversarial attacks, and manipulation techniques, allowing participants’ models to be tested under conditions that resemble those found in real-life scenarios, whether due to the rapid advancement in audio deepfake technologies, social media uploads, or unknown deepfake algorithms.
Furthermore, we provide a thorough analysis of the top-performing systems in each track in order to identify key strategies and common techniques employed by successful teams. This analysis, presented in Section 5, offers valuable insights into the current popular methodologies in the field of deepfake audio detection and analysis, and lays the groundwork for future research. Finally, we discuss the limitations observed in the current approaches and outline future directions for advancing research in deepfake audio detection and analysis in Section 6, highlighting the ongoing need for innovation and collaboration in this rapidly evolving area.
The primary contributions of this article include:
  1. (1)
    We detail the datasets used in all tracks of the ADD 2023 challenge and provide a comprehensive description of each of the datasets, which have been made publicly available for download;
  2. (2)
    We analyze the technical methodologies used by the top-performing participants in each task of the ADD 2023 challenge and note the commonalities and differences in their approaches;
  3. (3)
    We discuss the current technical limitations identified during the challenge and provide a roadmap for future research directions.
@@ 3 Evaluation Metrics
3.1 Track 1.1: Generation task (FG-G)
@@
3.2 Track 1.2: Detection task (FG-D)
@@
3.3 Track 2: Manipulation region location (RL)
The evaluation metric for the RL track is the sentence-level accuracy ($A_{s}$) and the segment-level $\text{F}_{1}$ score ($F_{1s}$). The overall performance of each team in the RL track (Score) is thus evaluated as a weighted sum of $A_{s}$ and the $F_{1s}$, with more emphasis on the latter, which is more challenging to achieve and also the main objective of this track, without ignoring the overall $A_{s}$, which is also important in real-world applications.
3.4 Track 3: Deepfake algorithm recognition (AR)
@@
4 Datasets
In this section, we present and describe the datasets [^2] http://addchallenge.cn/downloadADD2023 used in the ADD 2023 challenge. Each dataset is used for a specific task that corresponds to a track in the challenge.
In order to better represent real-world scenarios, the datasets are designed according to the following general principles: first, there should be a noticeable difference in the training and testing set data distributions, to ensure that models are able to generalize well to unseen data present in real-life applications; second, there needs to be a sense of attack-and-defence dynamic, which is especially important for the FG track; and third, the inclusion of unknown deepfake algorithms and manipulation techniques aims to further test the robustness of the participants’ systems.
4.1 Track 1.1: Generation task (FG-G)
The FG-G track being focused on the generation of deepfake audio, its dataset is composed of speech corpora with transcripts. In this case, we use the AISHELL-3 dataset [8] in the training process, which is a large-scale multi-speaker Mandarin speech dataset for TTS that contains roughly 85 hours of speech data from 218 native Mandarin speakers, totalling 88,035 utterances. This is done to ensure that the generated audio is of high quality and is able to emulate the characteristics of genuine audio.
In addition, the two-round setup of the FG-G sub-track and its dataset is designed to encourage participants to adapt their methods and models to new data and conditions, as well as to test the robustness of their systems. In this sense, the release of a benchmark model for the second round of the FG-G sub-track is intended to provide a target for participants to deceive, and to further motivate the development of more advanced and sophisticated deepfake audio generation methods.
The testing set, summarized in Table I, is composed of 998 text sentences (499 per each round) in Simplified Chinese characters, as well as four speaker IDs (two per each round, no overlap) from the aforementioned AISHELL-3 dataset. The average length of these sentences (excluding punctuations) is 66.98 characters, with a maximum length of 147 characters, a minimum length of 39 characters, and a variance of 136.78 characters.
@@ 4 Datasets
4.2 Track 1.2: Detection task (FG-D)
4.2.1 Training and development sets
The training and development sets of the FG-D track are deliberately designed to provide a more restricted set of data, reflecting the controlled and limited conditions under which detection models are typically trained and validated. This restriction is intentional, as it mirrors the reality that, during the training and validation phases, models often have access to a more homogenous and well-curated dataset. This dataset might lack the full diversity and complexity of the data that the model will encounter in practical, real-world applications.
In the training and development sets, only the AISHELL-3 dataset is used for genuine speech, totalling 60 speakers and 3,012 utterances for the training set and 60 speakers and 2,307 utterances for the development set (see Table II(a)). For fake audio, we use samples generated from the following methods:
4.2.2 Testing set
For the testing set of FG-D dataset, we include audio samples from a wider range of sources for both real and fake audio (see Table II(b)), since it is designed to be more challenging and diverse than the training and development sets, reflecting the broader range of scenarios that detection models are likely to encounter in real-world deployments, where the data may be more varied and unpredictable due to different environments, recording conditions, and deepfake manipulation techniques. The testing set, therefore, serves as a more challenging benchmark for evaluating the robustness and adaptability of the detection models.
The genuine audio portion of the Track 1.2 testing set comes from the following sources:
  • Tencent111111hhttps://www.tencentcloud.com/products/tts: A TTS and VC platform developed by Tencent providing high quality speech synthesis services.
By designing the dataset in this way, we seek to emphasize the importance of generalization, encouraging researchers to develop models that can perform well not just on the curated datasets they are trained on, but also on the unpredictable and diverse data they will face in actual use cases. This approach is crucial for advancing the field of deepfake audio detection, as it pushes the development of models that are not only accurate in controlled settings but also resilient and effective in the dynamic, real-world scenarios where they are most needed.
@@ 4.2.2 Testing set
4.3 Track 2: Manipulation region location (RL)
@@
As such, the setup of the RL track dataset aims to prompt the generalization of models to detect manipulation regions in real-world scenarios, where the data may be more varied and complex than in controlled settings. By providing a diverse and challenging dataset, we aim to push the development of models that can perform well not only on the data they are trained on but also on the unpredictable and diverse data they will encounter in practical applications.
4.3.1 Training and development sets
The training and development sets (see Table III(a)) contain genuine audio sourced from the training and development sets of the AISHELL-3 dataset and partially fake audio sourced from the training and development sets of the HAD dataset [28]. The choice of these datasets is motivated by the need to balance the complexity of the task with the availability of high-quality data. As the HAD dataset contains partially fake audio obtained by splicing genuine audio from AISHELL-3 with fake audio [28], it provides a more-or-less controlled environment for training and validating models to detect manipulation regions.
4.3.2 Testing set
The testing set is built to contain a more diverse and realistic set of scenarios than the training and development sets, reflecting the broader range of conditions that models are likely to face in real-world applications. By including a wider variety of sources and manipulation techniques, some of which may have been real audio spliced with other real audio, the testing set provides a more challenging benchmark for evaluating the performance of manipulation region location models.
The testing set of Track 2 (see Table III(b)) contains 1,030 speakers and 20,000 utterances for genuine audio from various sources like AISHELL-1, AISHELL-3, THCHS-30, Mobvoi, AliMeeting and Phone recordings. Despite the overlap in the sources of genuine audio between Track 1.2 and Track 2 testing sets, there is no overlap in actual utterances between the two tracks. The fake audio samples in the testing set are AISHELL-1 and AISHELL-2 samples spliced with either same-speaker recordings or fake audio generated with the Tacotron 2–LPCNet pipeline, with one or two spliced regions per sentence, as well as utterances processed by the Speech Edit model based on CampNet [25]. The fake audio samples were generated using the same speakers as the genuine audio samples. Genuine utterances are then spliced by replacing certain segments with fake audio or other recordings of the same speaker, following a similar procedure as in [28]. These partially-fake samples total 30,000 utterances from 148 speakers.
4.4 Track 3: Deepfake algorithm recognition (AR)
The objective of the AR track is to recognize the deepfake algorithm used to generate a given piece of audio. As such, the dataset for this track is designed to evaluate the ability of participants to recognize the deepfake algorithm. By including a diverse set of sources and manipulation techniques, both known and unknown, the dataset provides a benchmark for evaluating the performance of deepfake algorithm recognition models. The setup of the AR track dataset aims to prompt the generalization by the models to recognize deepfake algorithms in real-world scenarios, where the data may be more varied and complex than in controlled settings.
The dataset (see Table IV), is based on the datasets presented in [29, 30] is composed of genuine utterances from the AISHELL-3 dataset and fake utterances generated by various deepfake algorithms from the AISHELL-1 and AISHELL-3 text corpora. The deepfake algorithms used to generate the fake utterances include vocoders like HiFiGAN [9], WaveNet [14] and World [15], as well as commercial solutions from Aliyun [^13] https://ai.aliyun.com/nls/tts, DataBaker [^14] https://www.data-baker.com/specs/compose/online, and AISpeech [^15] https://beta.duiopen.com/openSource/technology/tts.
@@ 4.4 Track 3: Deepfake algorithm recognition (AR)
Further augmentations were made to the testing set to emulate real-world scenarios. With the aforementioned real and fake audio samples forming the “clean” condition, additional conditions were created to test the robustness of the deepfake algorithm recognition models. To simulate diverse acoustic environments of the real world, various types and levels of background noise were added to clean audio samples, creating the “noisy” condition to test deepfake algorithm recognition under suboptimal conditions. Additionally, to reflect audio compression commonly found on social media, the Jinshi Video Assistant software [^17] https://www.drmfab.cn/zhushou/ was used to compress clean audio, forming the “compressed” condition. These steps assess the algorithm’s adaptability to encoding changes.
TABLE V: DSR (%) and methods of top-performing systems in Track 1.1 (FG-G) submissions. [“Aug.” = augmentation; “Rep.” = representation; “spec” = spectrogram; “Arch.” = architecture; “AR” = auto-regressive]
@@
A01 [31] 44.97 Noise + reverb Phoneme seq. Mel spec Mel spec Tacotron 2 WaveRNN - BiLSTM AR
A02 [32] 43.63 - Phoneme seq. Mel spec Mel spec FastPitch HiFiGAN MFA One-hot NAR
A03 [33] 41.48 - Context encoder Linear spec Waveform Hier-TTS - - One-hot AR
A05 [34] 37.35 Concatenation G2P+BERT Mel spec Mel spec FastSpeech 2 HiFiGAN Kaldi One-hot NAR
@@
A06 [35] 30.69 - Phoneme seq. Linear spec Waveform VITS - MAS One-hot NAR
@@
5 Technical Analysis
We analyze the technical details of the top-5 performing teams of each track in this section, in order to provide insights into the strategies and techniques that led to their success and to identify common trends and best practices in the field of deepfake audio detection and analysis. In the tables are presented not only the performances of top-performing participating teams, but also the results of the baseline systems [6] (denoted in the form of S#, where # is the baseline number). Among the baselines, S01S03 are the baseline systems for the FG-D sub-track, S04 is the baseline systems for the RL track, and S05S06 are the baseline systems for the AR track.
5.1 Track 1.1: Generation task (FG-G)
@@ 5 Technical Analysis
    Data augmentation: The use of data augmentation in Track 1.1 is somewhat limited compared to other tracks, and the teams that adopted it, A01 [31] and A05 [34], used it in the process of speaker ID and characteristic modelling rather than in the audio generation process. Of those two teams, A01 uses reverberation and additive noises to augment the data, while A05 concatenates and re-splices training data belonging to the same speaker ID. This might be due to the fact that the AISHELL-3 dataset is already quite large, and the use of data augmentation might not be as crucial as in other tasks.
@@
    Text representation: Most teams opted for phoneme sequence as the text representation, which is a common choice for TTS systems, as it is seen to be closer to the phonetic representation of speech, and thus more suitable for speech synthesis, as well as being more robust to spelling errors and out-of-vocabulary words. A notable exception to this trend is Team A03 [33], which used a context encoder to encode the input text into a fixed-length vector. Another deviation from this is Team A05 [34], who used a grapheme-to-phoneme model to convert the input text into phoneme sequence, and used BERT for disambiguating homographs. This approach is more robust to spelling errors and out-of-vocabulary words, as it can handle unseen words by converting them into phonemic representations which are more directly related to the speech signal and also present in the training data in some form.
@@
    Architectures: The use of non-autoregressive models like FastSpeech 2 is common in this track, which allows for variations in generation pipelines in the form of duration models. Aside from using models that are non-autoregressive, most teams adopt approaches with intermediate audio representations like Mel spectrogram, which are then passed through vocoder models to generate the final audio waveform. It is worth noting, however, that A06 [35] used the VITS architecture, which is fully end-to-end, with its own alignment module of Monotonic Alignment Search, that allows for the generation of audio waveform directly from phoneme sequence by using a variational autoencoder-like encoder-decoder architecture to directly decode into audio waveform. This approach is more computationally efficient and allows for faster inference times, as it does not require the intermediate step of generating Mel spectrograms.
TABLE VII: Scores (%), sentence accuracies $A_{s}$ (%), segment-wise F1 scores (%) and methods of top-performing systems and baseline in Track 2 (RL). [“Aug.” = augmentation]
Team Score ($\uparrow$) $A_{s}$ ($\uparrow$) $F_{1s}$ ($\uparrow$) Data Aug. Acoustic Features Back-end Classifiers Model Fusion
C01 [42] 67.13 82.23 60.66 Noise; reverb Wav2Vec 2.0 ResNet–Transformer–LSTM -
@@
C05 [46] 59.12 74.52 52.53 Noise Raw, Wav2Vec 2.0 AASIST, FC layer Weighted segment-wise
@@
5.2 Track 1.2: Detection task (FG-D)
@@
    Data augmentation: The use of data augmentation on the training data in Track 1.2 is widespread and has found much success, with most teams opting to use MUSAN [52] noise augmentation. The use of reverberation is also popular, with most teams opting to add reverberation to their training data to improve the robustness of their models. This is likely due to the fact that the testing set contains a wide variety of both real and fake audio samples with different conditions, and thus data augmentation may be crucial to ensure that the model is robust to these different conditions. Interestingly, Team B01 [36] opted to furthermore use copy synthesis as one of their data augmentation methods to obtain more training data from the AISHELL-3 corpus, and given the success of their system, this approach may have been effective in the task.
@@
5.3 Track 2: Manipulation region location (RL)